# Convergence Acceleration of Favre-Averaged Non-Linear Harmonic Method

### v1

@@ Introduction
Generalized RK Scheme for FNLH
The above RK scheme will be applied both to the mean flow and the time-linearized harmonic flow. For the mean flow, $\Gamma$ and $\mathbf{K}$ can be written as [25]:
where $\mathbf{J}$ is the flux Jacobian matrix $\frac{\partial{\mathbf{R}}}{\partial\mathbf{Q}}$. The flux Jacobian is assembled using the information of the immediate neighbors with first-order spatial accuracy, and hence it is an approximation of the true Jacobian. $\mathbf{P}$ is the diagonal block of the Jacobian matrix $\mathbf{J}$. For the explicit scheme, $\Gamma$ is the the Courant–Friedrichs–Levy (CFL) number. For the implicit scheme, $\Gamma$ is related to the relaxation factor $\varepsilon$ and in this work a value of $0.6$ [31] is used.
@@
With respect to the time-linearized flow, the contribution of the spectral source term must be taken into account, and consequently a diagonal matrix $I\omega_{l}\mathbf{I}\mathbf{V}$ is added to $\mathbf{K}$, where $\omega_{l}$ is the angular frequency. This leads to the preconditioner matrix for each harmonic $\mathbf{\hat{K}}_{l}$, and this is shown in Equation 10.
@@
$\mathbf{\hat{K}}_{l}=\left\{\begin{array}[]{ll}\mathbf{P}+I\omega_{l}\mathbf{I}\mathbf{V}&\text{Explicit}\\ \frac{\mathbf{P}}{\varepsilon\sigma_{\text{CFL}}}+\mathbf{J}+I\omega_{l}\mathbf{I}\mathbf{V}&\text{Implicit}\end{array}\right.$ (10)
@@
in which $I$ is the imaginary unit and $\mathbf{I}$ is the unit matrix. Since $\omega_{l}$ is generally different for each harmonic, $\mathbf{\hat{K}}_{l}$ also differs from harmonic to harmonic. It is not practical to store the LHS for each harmonic separately, as this will make the scheme not memory efficient. Since the sparse pattern of the matrix remains the same for each harmonic, this feature can be explored to develop a memory-efficient approach to solving FNLH.
@@
For each RK stage (either for the mean flow or the time-linearized harmonic flow), an update of the solution $\Delta\mathbf{Q}^{(k)}$ to the vector of the intermediary solution $\mathbf{Q}^{n,(k)}$ must be calculated, and this reads as follows:
$\Delta\mathbf{Q}^{(k)}=\Gamma\alpha^{(k)}\mathbf{K}^{-1}(\mathbf{R}^{(k)}_{i}-\mathbf{R}^{*})$ (11)
@@
To use more traditional symbols for the discussion of solving linear systems, Equation 11 can be written in a more familiar form. For example, the one for the mean flow can be written as $\mathbf{A}\mathbf{x}=\mathbf{B}$, where $\mathbf{A}=(\Gamma\alpha^{(k)}\mathbf{K}^{-1})^{-1}$, $\mathbf{B}=(\mathbf{R}^{(k-1)}_{i}-\mathbf{R}^{*}_{i})$ and $\mathbf{x}$ is $\Delta\mathbf{Q}^{(k)}_{i}$. For the explicit scheme, $\mathbf{A}$ is a block diagonal matrix, and its inversion can be performed point by point. This is essentially a block Jacobi method, and inverting the block diagonal matrix is a simple and fast process. For the implicit scheme, the LHS of the mean flow or the time-linearized harmonic flows is a large sparse matrix. Applying an exact inversion of this matrix is seldom economical, because the improvement of the convergence rate is not good enough to justify the high computational cost of inverting this matrix exactly. In this work, the procedure similar to that proposed by Swanson et al. [22] is followed. This matrix is inexactly inverted and the linear system is solved iteratively. The incomplete lower-upper decomposition with zero fill (ILU(0)) is used to approximate the matrix inversion.
@@
With respect to the time-linearized harmonic flow, after inspecting Equation 10 and Equation 9, it can be seen that the LHS of the linearized flow has the same sparse pattern as the one for the nonlinear flow. The difference comes mainly from the diagonal block. Furthermore, if there are periodic boundaries in the simulation, phase-lagged boundary conditions are applied on the periodic boundaries. Since the interblade phase angle will be different for each harmonic, this will also lead to a difference of the LHS for each harmonic, but it has no impact on the sparse pattern of the LHS. Based on the above observation, the coupled system of FNLH for each nonlinear iteration can be symbolically represented as:
$\mathbf{A}\mathbf{x}$ $=\mathbf{B}$ mean flow
$(\mathbf{A}+I\omega_{l}\mathbf{I}\mathbf{V})\hat{\mathbf{x}}_{l}$ $=\hat{\mathbf{B}}_{l}$ linearized flow (12)
@@
in which $\mathbf{A}$ represents the LHS of the non-linear flow. $\mathbf{B}$ and $\mathbf{\hat{B}}_{l}$ denote the RHS of the mean flow and the RHS of the $l^{\text{th}}$ harmonic.
@@
Regarding the coherence of the sparse pattern of the preconditioner matrix from the nonlinear flow and the linearized flow, the flux Jacobian matrix $\mathbf{J}$ only needs to be assembled once by the nonlinear flow for each nonlinear iteration, and is then reused by the linearized flow. From an implementation point of view, the mean flow and the first harmonic of the linearized flow will allocate memory for its LHS, respectively, but the LHS of all the harmonics of the linearized flow reuses the memory space of that of the first harmonic. This treatment is used for both the explicit and implicit schemes and allows the memory usage of FNLH to not depend on the number of harmonics to compute; therefore, the resulting procedure is memory efficient.
Number Representation
@@
The sparse linear system from the time-linearized flow consists of complex numbers. We can either cast this complex linear system into its equivalent system with real numbers, which is shown in Equation 13, or directly solve this complex linear system.
@@
$\begin{bmatrix}\mathbf{A}&-{\omega}_{l}\mathbf{I}\mathbf{V}\\ {\omega}_{l}\mathbf{I}\mathbf{V}&\mathbf{A}\end{bmatrix}\begin{bmatrix}\text{Re}\{\Delta\hat{\mathbf{x}}_{l}\}\\ \text{Im}\{\Delta\hat{\mathbf{x}}_{l}\}\end{bmatrix}=\begin{bmatrix}\text{Re}\{\hat{\mathbf{B}}_{l}\}\\ \text{Im}\{\hat{\mathbf{B}}_{l}\}\end{bmatrix}$ (13)
@@
The benefit of using real numbers is that existing implementation of incomplete matrix factorization can be reused. However, as shown in Equation 13, the sparse pattern of the new LHS differs from that of the mean flow solver and this means that a new lookup table is required for the matrix factorization (i.e. ILU(0)). For the approach using complex numbers, the sparse pattern of the LHS is the same as that of the mean flow solver, and this also facilitates parallel implementation (i.e. parallel domain decomposition) of incomplete matrix factorization. In the authors’ opinion, these benefits outweigh the effort to implement an incomplete matrix factorization with complex numbers. Therefore, in this work, the approach with complex numbers is chosen to solve the sparse linear system of the time-linearized flow.
@@
For parallel implementation of the implicit scheme, a domain decomposition strategy is required to partition the LHS of the mean flow or the linearized flow for different processors. As the sparse patterns of the linearized flow and the mean flow are identical, the same domain decomposition strategy is used. In this work, the partition is done in such a way that the entries related to two nodes that are in different partitions are ignored. This can be demonstrated in Fig. 3, which shows the sparse pattern and the partition of the LHS on four processors. Each of the 4 boxes encloses a sub-linear system for each processor. These sub-linear systems only couple nodes that are partitioned into the same processor. This effectively leads to a parallelisable set of subproblems for each processor. Furthermore, as was discussed in the previous section, the linear system with complex numbers is solved for the linearized flow. If an equivalent linear system with real numbers is solved, the off-diagonal part, which contains the contributions of the spectral source term, could be ignored for parallel computations. This could lead to potentially numerical problems.
In addition, as the coupling of the nodes that are in different processors is ignored, the convergence rate of the implicit solver will degrade as more processors are used. For the extreme case, where the number of processors is equal to the number of nodes, the parallel implicit solver will degrade to the explicit scheme. It is found that serious convergence degradation can occur if the total number of nodes to the number of processors is less than $O(10^{3})$ [31]. On the other hand, in reality, to obtain good scalability for parallel computation, this ratio normally is at least $O(10^{4})$. Therefore, in practice, the convergence rate will only have a weak dependence on the number of processors, and this effect will be demonstrated in the test cases.
Figure 3: Parallel domain decomposition for the LHS matrix.
@@ Results
We demonstrate the computational performance of the explicit and implicit versions of the generalized RK scheme for FNLH and perform a comparative study of both approaches in terms of convergence, wall clock time and memory consumption. Two test cases are used to serve this purpose. The first case is to model the rotor-rotor interactions in a compressor, and the second case is to trace hot streak migrations in turbine stages. Regarding verification and validation, the performance of FNLH for these two cases has been demonstrated in the previous work [4] for multirow coupling. Nevertheless, comparisons with URANS are still provided to demonstrate that the explicit/implicit scheme converges to the correct results. Since FNLH can be considered as a reduced model of URANS, comparisons with URANS suffice for validation. Besides, the capability of the bladerow interface in HYDRA to correctly transmit disturbances in a multi-row environment is also verified by comparing with analytical solutions.
The key parameters of the generalized RK scheme are summarized in Table 2. For the implicit solver, the incomplete LU factorization with zero fill-in level is used and is denoted as ILU(0). A fixed threshold of 0.01 is used for the linear residual to iteratively solve the sparse linear system for both the mean flow and the time-linearized harmonic flow. The maximum number of iterations to solve the sparse linear systems is 10. For the explicit scheme, a V cycle is used for the multigrid and the order of spatial accuracy for the coarse grid levels is set to 1.
@@
Verification of Multirow Bladerow Interface in HYDRA
The effectiveness of the bladerow interface in transmitting disturbances in a multiple bladerow setting is demonstrated here. Figure 4 a) shows the configuration of three ducts that are connected to each other in the streamwise direction. ”duct2” is rotating in the direction indicated in the figure. ”duct1” and ”duct 3” are stationary. The pitch of the ducts are $\frac{1}{36}\pi$, $\frac{1}{24}\pi$, and $\frac{1}{36}\pi$, respectively. The flow is assumed to be inviscid, and a total temperature distortion is specified at the inlet of ”duct1”. The magnitude of the distortion is $5\%$ of the inlet total temperature. As there is no work input or heat exchange, theoretically the total temperature distortion should be transported downstream by the background flow without attenuation.
Figure 4: Transmitting a total temperature distortion through 3 ducts.
Structured meshes are used for all the three ducts and two sets of grid resolution are employed. The first mesh has a resolution of $60\times 150\times 1$ and the second has $100\times 200\times 1$. They represent resolutions in the streamwise, circumferential, and spanwise directions, respectively. Two background flows are used. For the first, the flow is axial, and for the second, the whirl angle is $45^{\circ}$. This is shown in Fig. 4 b) and c). The flow condition (e.g. reduced frequency) is representative of typical turbomachinery flow conditions. The total temperature distribution in any downstream position can be obtained analytically by applying the phase shift of the inlet total temperature distribution if the background flow is non-axial. Figure 5 shows the comparisons between the FNLH solution and the analytical solution. There is good agreement between FNLH and analytical solutions. As the mesh is refined, the agreement is improved due to reduced numerical dissipation.
Figure 5: Comparisons of total temperature distribution between FNLH and analytical solution.
@@
This test case is the mid-span stream tube of the front 1.5 stages of an 8-stage high-speed machine [32], and the stream tube accounts for roughly $4.8\%$ of the span height. The first- and second-stage rotors of this 8-stage compressor are used [4]. Blade counts are scaled to the ratio of 3(R1):5(S1):4(R2) to reduce the computational cost of URANS. The in-house meshing tool [33] is used to create multiblock structured meshes to discretize the computational domain in the blade-to-blade section and will be used to mesh the computational domain for all test cases. In the spanwise direction, one layer of hexahedral mesh is used and the total hexahedral element is 33090. The average $y^{+}$ on the blade surface is around 5. The Sparlart-Allmaras turbulence model [34] is used by default in all test cases. Non-reflective treatment for mean flow [35] and linearized flow [30] is applied at the bladerow interfaces. For the URANS simulation, single-passage meshes are replicated in the circumferential direction to form a $30^{\circ}$ sector. 500 time steps are used for this $30^{\circ}$ sector simulation, which corresponds to 100 time steps for the stator to sweep one passage of the rotor that has the smaller pitch.
Table 3 shows the FNLH setup for this compressor case, and $\Omega$ is the shaft speed. 15 harmonics are used in total. In the current FNLH implementation in Hydra, all bladerows compute the same total number of harmonics. For R1, 15 harmonics are used to calculate the potential fields of S1. For S1, 12 harmonics are used to model the R1 wake and 3 harmonics are used to compute the potential field of R2. For R2, 8 harmonics are used for the wake of S1 and 5 harmonics are used to model the wake of R1. For the harmonic representation of the R1 wake in R2, its wave number is a linear combination of the blade counts of R1 and S1, and its temporal frequency is related to the shaft speed $\Omega$. Such an FNLH setup was used in our previous work, and for more details, the readers can refer to the previous work [4].
@@
R1 $i\text{N}_{S1}\Omega,i\in[1:15]$ $i\text{N}_{S1},i\in[1:15]$ exit, S1 potential field
S1 $i\text{N}_{R1}\Omega,i\in[1:12]$ $i\text{N}_{R1},i\in[1:12]$ inlet, R1 wake
$-i\text{N}_{R2}\Omega,i\in[1:3]$ $i\text{N}_{R2},i\in[1:3]$ exit, R2 potential field
R2 $i\text{N}_{S1}\Omega,i\in[1:8]$ $i\text{N}_{S2},i\in[1:8]$ inlet, S1 wake
@@
Figure 6 shows the unsteady entropy fields for this 1.5 stage transonic compressor computed by FNLH and URANS. Entropy is evaluated as $s=\frac{R}{\gamma-1}(\ln(p)-\gamma\ln(\rho))$. FNLH shows good agreement with URANS, in particular, the wake of R1 is reconstructed satisfactorily in R2. Figure 7 shows the circumferential distribution of the entropy at two streamwise locations, which are located at the midway betweeen S1 inlet and LE, R2 inlet and LE, respectively. Good agreement with URANS data is observed. The results are computed by the implicit RK scheme, hence we are confident that the generalized RK scheme still produces the correct results.
@@
Figure 8 shows the RMS of the residuals for the mean flow and the harmonic flows from the implicit scheme. For example, the residual for ”Harmonic-0” is the residual of the 1st harmonic computed by all the blade rows. According to the FNLH setup in Table 3, Harmonic [0-7] represents adjacent row interactions for all the blade rows. Harmonic [8:14] still computes adjacent row interaction for R1 and S1, but for R2 they model R1 wake in the computational domain of R2. For adjacent row interactions, harmonics with larger indices have shorter wave lengths and higher frequencies; Fig. 8 shows that these harmonics tend to have faster convergence. This can be explained by the fact that, because of the shorter wave length, the wave length to grid size ratio is smaller, and this means that these harmonics are more dissipated by the numerical scheme. Furthermore, as the spectral source term only contributes to the diagonal part of the LHS of the linearized flow (Equation 10), this will enhance the diagonal dominance and consequently improve the convergence rate.
Figure 8: Convergence history of all the harmonics with the implicit scheme for the 1.5 stage compressor case.
Since a dozen harmonics could be computed in a FNLH run, it is not convenient to inspect the residuals for each harmonic. Therefore, the RMS of the residuals for all the harmonics is monitored. This is defined as:

### v2

@@ Introduction
Generalized RK Scheme for Time-Linearized Harmonic Flows
With respect to the time-linearized flow, the structure of the generalized RK scheme for the mean flow can be re-used. $\mathbf{Q}^{n,(k)}$ should be replaced with the harmonic flow perturbations $\mathbf{\hat{Q}}^{n,(k)}_{l}$, and $\mathbf{R}^{(k)}$ should be replaced with the non-linear residuals from the time-linearized flows $\mathbf{\hat{R}}^{(k)}_{l}$, for example:
$\mathbf{\hat{Q}}^{n,(0)}_{l}$ $=\mathbf{\hat{Q}}^{n}_{l}$
$\mathbf{\hat{Q}}^{n,(k)}_{l}$ $=\mathbf{\hat{Q}}^{n,(0)}-\Gamma\alpha^{(k)}\mathbf{\hat{K}}^{-1}_{l}(\mathbf{\hat{R}}^{(k)}_{l}-\mathbf{\hat{R}}^{*}_{l}),k=[1:5]$
$\mathbf{\hat{Q}}^{n+1}_{l}$ $=\mathbf{\hat{Q}}^{n,(5)}_{l}$ (12)
@@
in which $\Gamma$ is the same as the mean flow RK scheme, some attention is required for the matrix preconditioner. Due to the spectral source term in the time-linearized solver, a diagonal matrix $I\omega_{l}\mathbf{I}V$ is added to the diagonal block of $\mathbf{K}$, where $\omega_{l}$ is the angular frequency. This leads to the preconditioner matrix for each harmonic $\mathbf{\hat{K}}_{l}$, which consists of complex numbers, and this is shown in Equation 13.
@@
$\mathbf{\hat{K}}_{l}=\left\{\begin{array}[]{ll}\mathbf{P}+I\omega_{l}\mathbf{I}V&\text{Explicit}\\ \frac{\mathbf{P}}{\varepsilon\sigma_{\text{CFL}}}+\mathbf{J}+I\omega_{l}\mathbf{I}V&\text{Implicit}\end{array}\right.$ (13)
@@
in which $I$ is the imaginary unit, $V$ is the volume of a control volume, and $\mathbf{I}$ is the unit matrix whose size is $5\times 5$ for a 3D problem. Because $\omega_{l}$ and the interblade phase angle are generally different for each harmonic, $\mathbf{\hat{K}}_{l}$ is then different from harmonic to harmonic, but the sparse pattern of $\mathbf{\hat{K}}_{l}$ remains the same as the mean flow. This feature can be explored to develop a memory-efficient approach to solving FNLH.
Solution Update
@@
For each RK stage of the mean flow, an update of the solution $\Delta\mathbf{Q}^{(k)}$ to the vector of the intermediary solution $\mathbf{Q}^{n,(k)}$ must be calculated, and this reads as follows:
$\Delta\mathbf{Q}^{(k)}=\mathbf{Q}^{n,(k)}-\mathbf{Q}^{n,(0)}=-\Gamma\alpha^{(k)}\mathbf{K}^{-1}(\mathbf{R}^{(k)}_{i}-\mathbf{R}^{*})$ (14)
To use more traditional symbols for the discussion of solving sparse linear systems, Equation 14 can be written in a more familiar form $\mathbf{A}\mathbf{x}=\mathbf{B}$. In the following, $\mathbf{A}$ refers to the left-hand side (LHS) of the sparse linear system and $\mathbf{B}$ refers to the right-hand side of this system. For the mean flow, they represent:
$\mathbf{A}$ $=(\Gamma\alpha^{(k)}\mathbf{K}^{-1})^{-1}=\frac{\mathbf{K}}{\Gamma\alpha^{k}}$ (15)
$\mathbf{B}$ $=-(\mathbf{R}^{(k)}-\mathbf{R}^{*})$ (16)
$\mathbf{x}$ $=\Delta\mathbf{Q}^{(k)}$ (17)
@@
For the explicit scheme, $\mathbf{A}$ is a block diagonal matrix, and its inversion can be performed point by point. This is essentially a block Jacobi method, and inverting the block-diagonal matrix is a simple and fast process. For the implicit scheme, $\mathbf{A}$ is a large sparse matrix. Solving the linear system exactly is seldom economical, because the improvement of the convergence rate is not good enough to justify the high computational cost of inverting this matrix exactly. In this work, the procedure similar to that proposed by Swanson et al. [22] is followed. This matrix is inexactly inverted, and the linear system is solved iteratively. The incomplete lower-upper decomposition with zero level of fill-in (which is commonly denoted as ILU(0)) is used to approximate the matrix inversion. To be more specific, if an ILU(0) procedure is applied to the matrix $\mathbf{A}$. i.e. $\mathbf{A}\approx\mathbf{L}\mathbf{U}$. $\mathbf{L}$ is a lower triangular matrix and $\mathbf{U}$ is an upper triangular matrix. By decomposing the original matrix to a lower and upper triangular matrix, the matrix inversion can be computed efficiently. The zero level of fill-in means that the sparsity pattern of $\mathbf{L}$ and $\mathbf{U}$ is chosen to be the same as the sparsity pattern of the matrix $\mathbf{A}$, which means if an entry in $\mathbf{A}$ is zero, then that entry in the matrix of $\mathbf{L}$ or $\mathbf{U}$ is also zero. Therefore, $\mathbf{L}\mathbf{U}$ is just an approximation of $\mathbf{A}$, and so will the matrix inversion by ILU(0). Because there is no need to solve the linear system exactly, ILU(0) is a popular choice for matrix inversion to solve the linear system.
@@
With respect to the time-linearized harmonic flow, after inspecting Equation 13 and Equation 8, it can be seen that the LHS of the linear system of the linearized flow has the same sparse pattern as the one for the nonlinear flow. The difference comes mainly from the diagonal block. Furthermore, if there are periodic boundaries in the simulation, phase-lagged boundary conditions are applied on the periodic boundaries. Since the interblade phase angle will be different for each harmonic, this will also lead to a difference of the LHS for each harmonic, but it has no impact on the sparse pattern of the LHS. The linear system for the $l^{th}$ harmonic can be written as:
$\frac{\mathbf{K}_{l}}{\Gamma\alpha^{k}}\hat{\mathbf{x}}_{l}=\hat{\mathbf{B}}_{l}$ (18)
@@
in which $\mathbf{\hat{B}}_{l}$ and $\mathbf{\hat{x}}_{l}$ denote the RHS and solution update of the linear system of the $l^{\text{th}}$ harmonic, respectively. Furthermore, the LHS of this system can be written using that of the mean flow as:
$(\mathbf{A}+\frac{I\omega_{l}\mathbf{I}V}{\Gamma\alpha^{k}})\hat{\mathbf{x}}_{l}=\hat{\mathbf{B}}_{l}$ (19)
@@
When comparing the LHS of the linear system of the mean flow (Equation 15) and that of the linearized flow(Equation 19), it can be seen that they share the same sparse pattern. This means that the residual Jacobian matrix $\mathbf{J}$ only needs to be assembled once by the nonlinear flow for each nonlinear iteration, and is then reused by the linearized flow. From an implementation point of view, the mean flow and the first harmonic of the linearized flow will allocate memory for its LHS, respectively, but the LHS of all the harmonics of the linearized flow reuses the memory space of that of the first harmonic. This treatment is used for both explicit and implicit schemes and allows the memory usage of FNLH to not depend on the number of harmonics to compute; therefore, the resulting procedure is memory efficient. It is worth mentioning that a similar procedure was also explored in the harmonic balance method [13], however, this procedure is described here in the context of FNLH.
Complex Number Representation
@@
The sparse linear system of the time-linearized flow consists of complex numbers. We can cast this complex linear system (i.e. Equation 19) into its equivalent system with real numbers, or we can directly solve this complex linear system.
@@
$\begin{bmatrix}\mathbf{A}&-{\omega}_{l}\mathbf{I}\mathbf{V}/(\Gamma\alpha^{k})\\ {\omega}_{l}\mathbf{I}\mathbf{V}/(\Gamma\alpha^{k})&\mathbf{A}\end{bmatrix}\begin{bmatrix}\text{Re}\{\Delta\hat{\mathbf{x}}_{l}\}\\ \text{Im}\{\Delta\hat{\mathbf{x}}_{l}\}\end{bmatrix}=\begin{bmatrix}\text{Re}\{\hat{\mathbf{B}}_{l}\}\\ \text{Im}\{\hat{\mathbf{B}}_{l}\}\end{bmatrix}$ (20)
@@
The benefit of using real numbers is that the existing implementation of incomplete matrix factorization can be reused. However, as shown in Equation 20, the sparse pattern of the new LHS differs from that of the mean flow solver and this means that a new lookup table is required for the matrix factorization (i.e. ILU(0)). For the approach using complex numbers, the sparse pattern of the LHS is the same as that of the mean flow solver, and this also facilitates parallel implementation (i.e. parallel domain decomposition) of incomplete matrix factorization. However, a complex-number version of the baseline ILU(0) procedure needs to be implemented. From the authors’ point of view, turning an incomplete matrix factorization into its complex version is much easier than implementing a new look-up table and its related domain decomposition. Therefore, in this work, the approach with complex numbers is chosen.
@@
For parallel implementation of the implicit scheme for both the mean flow and the harmonic flows, a domain decomposition strategy is required to partition the LHS of the sparse linear system of the mean flow (Equation 15) or the linearized flow (Equation 19) for different processors. As the sparse patterns of the linearized flow and the mean flow are identical, the same domain decomposition strategy is used. In this work, the partition is done in such a way that the entries related to two nodes that are in different partitions are ignored. This can be demonstrated in Fig. 4. It shows the sparse pattern of the LHS for the mean flow of a compressor blade passage that is meshed with a multi-block structured grid, and its partition of the LHS on four processors. Each of the 4 boxes encloses a sub-linear system for each processor. These sub-linear systems only couple nodes that are partitioned into the same processor. This effectively leads to a parallelisable set of sub-problems for each processor.
In addition, as the coupling of the nodes that are in different processors is ignored, the convergence rate of the implicit solver will degrade as more processors are used. For the extreme case, where the number of processors is equal to the number of nodes, the parallel implicit solver will degrade to the explicit scheme. It is found that serious convergence degradation can occur if the total number of nodes to the number of processors is less than $O(10^{3})$ [32]. On the other hand, to obtain a good scalability for parallel computation, this ratio should normally be at least $O(10^{4})$. Therefore, in practice, the convergence rate will only have a weak dependence on the number of processors, and this effect will be demonstrated in the test cases.
Figure 4: Parallel domain decomposition for incomplete matrix factorization.
@@ Results
We demonstrate the computational performance of the explicit and implicit versions of the generalized RK scheme for FNLH and perform a comparative study of both approaches in terms of convergence, wall clock time and memory consumption. Three test cases are used to serve this purpose. The first case is to model the rotor-rotor interactions in a compressor, the second case is to trace hot streak migrations in turbine stages, and the third test case is the low-pressure compression system of a turbofan engine.
The key parameters of the generalized RK scheme are summarized in Table 2. For the implicit solver, the incomplete LU factorization with zero fill-in level is used and is denoted as ILU(0). A fixed threshold of 0.01 is used for the linear residual to iteratively solve the sparse linear system for both the mean flow and the time-linearized harmonic flow. Taking the mean flow as an example, 0.01 means that $||\mathbf{B}-\mathbf{A}\mathbf{x}||_{2}$ drops by two orders of magnitude compared to its initial value. In addition, the maximum number of iterations to solve the sparse linear systems is 10. For the explicit scheme, a V cycle is used for the multigrid and the order of spatial accuracy for the coarse grid levels is set to 1.
@@
This test case is the mid-span stream tube of the front 1.5 stages of an 8-stage high-speed machine [34], and the stream tube accounts for approximately $4.8\%$ of the span height. The first- and second-stage rotors of this 8-stage compressor are used [4]. Blade counts are scaled to the ratio of 3(R1):5(S1):4(R2) to reduce the computational cost of URANS. The in-house meshing tool [35] is used to create multiblock structured meshes to discretize the computational domain in the blade-to-blade section and will be used to mesh the computational domain for all test cases. In the spanwise direction, one layer of hexahedral mesh is used, and the total hexahedral element is 33090. The average $y^{+}$ on the blade surface is around 5. The Sparlart-Allmaras turbulence model [36] is used by default in all test cases. Non-reflective treatment for mean flow [37] and linearized flow [31] is applied at the bladerow interfaces. For the URANS simulation, single-passage meshes are replicated in the circumferential direction to form a $30^{\circ}$ sector. Dual-time stepping is used and 500 time steps are used for this simulation of the $30^{\circ}$ sector, which corresponds to 100 time steps for the stator to sweep one passage of the rotor that has the smaller pitch.
Table 3 shows the FNLH setup for this compressor case, and $\Omega$ is the shaft speed. 15 harmonics are used in total. In the current FNLH implementation in Hydra, all bladerows compute the same total number of harmonics. For R1, 15 harmonics are used to calculate the potential fields of S1. In the “note” column, it shows that these modes are applied on the exit boundary of R1 and are used to calculate the S1 potential field. For S1, 12 harmonics are used to model the R1 wake and they are applied at the inlet of S1. 3 harmonics are used to compute the potential field of R2 and are applied at the exit boundary of S1. This information is also provided in the “note ” column. For R2, 8 harmonics are used for the wake of S1 and 7 harmonics are used to model the wake of R1. For the harmonic representation of the R1 wake in R2, its wave number is a linear combination of the blade counts of R1 and S1, and its temporal frequency is related to the shaft speed $\Omega$. These harmonics are organized into three groups based on the value of $m$ and are referred to as “wake-1”, “wake-2” and “wake-3”, respectively. This FNLH setup was used in our previous work, and for more details, the readers can refer to the previous work [4].
@@
R1 $i\text{N}_{S1}\Omega,i\in\{1,2,...15\}$ $i\text{N}_{S1},i\in\{1,2,...15\}$ exit, S1 potential field
S1 $i\text{N}_{R1}\Omega,i\in\{1,2,...12\}$ $i\text{N}_{R1},i\in\{1,2,...12\}$ inlet, R1 wake
$-i\text{N}_{R2}\Omega,i\in\{1,2,3\}$ $i\text{N}_{R2},i\in\{1,2,3\}$ exit, R2 potential field
R2 $i\text{N}_{S1}\Omega,i\in\{1,2,...8\}$ $i\text{N}_{S2},i\in\{1,2,...8\}$ inlet, S1 wake
@@
Figure 5 shows the instantaneous entropy fields for this 1.5-stage transonic compressor computed by FNLH and URANS. The entropy is normalized as ($S-S_{min})/(S_{max}-S_{min})$. $S_{min}$ and $S_{max}$ are the minimum and maximum entropy in the circumferential direction at a plane that is midway between the S1 inlet and the S1 LE. FNLH shows good agreement with URANS, in particular, the wake of R1 is satisfactorily reconstructed in R2. Figure 6 shows the circumferential distributions of the normalized entropy at two streamwise locations, which are located midway between the S1 inlet and LE, the R2 inlet and LE, respectively. Good agreement with URANS data is observed. The results are computed by the implicit RK scheme, and hence we are confident that the generalized RK scheme still produces the correct results. Nevertheless, some discrepancies between FNLH and URANS data can still be observed; this is because frozen turbulence is assumed and cross-coupling is not included; future work is to address these issues and enhance the accuracy of the Hydra FNLH solver.
@@
Figure 7 shows the RMS of the residuals for the mean flow and the harmonic flows from the explicit and implicit schemes. For example, the residual for ”Harmonic-0” is the residual of the first harmonic computed by all the blade rows. According to the FNLH setup in Table 3, Harmonic [0-7] represents the adjacent row interactions for all the blade rows. Harmonic [8-14] still computes the adjacent row interaction for R1 and S1, but for R2 they model the R1 wake in the computational domain of R2. Figure 7 shows that for both implicit and explicit schemes, the harmonics converge at difference rates, and the harmonic that has a lower convergence rate in an explicit scheme also generally has a lower rate in the implicit scheme. In particular, for adjacent row interactions, harmonics with larger indices have shorter wave lengths and higher frequencies, and Fig. 7 shows that these harmonics tend to have faster convergence. This could be explained by the fact that, because of the shorter wave length, the wave length to grid size ratio is smaller, and this means that these harmonics are more dissipated by the numerical scheme.
Figure 7: Convergence history of all the harmonics with the implicit and explicit schemes for the 1.5-stage compressor case.
Since a dozen harmonics could be computed in an FNLH run, it is not convenient to inspect the residuals for each harmonic. Therefore, the RMS of the residuals for all harmonics is monitored. This is defined as: